15 January 2008

RHIC Running:  Yesterday's attempts to raise chromaticity by one unit gave somewhat better Yellow lifetimes and reduced coherence in the beams, but lower overall luminosity.  Then tunes at transition were increased slightly to compensate for the tune shift observed across the bunches due apparently to electron cloud.  This gave somewhat more luminosity. 

RF would like to put in the circuit for quad mode damping.  It can be put in behind the store, but it should be done at a time when no other changes are made so that any effects of the damping can be easily determined.  This will be done for the new fill at 13:00 today.

The AGS Cold Snake is now cold.  It has been hi-opt tested, and will get a full power test tomorrow.


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020